Find the height and base of the parallelogram that has an area of 100 square units.
Height=x
base=x+15
: Find the height and base of the parallelogram that has an area of 100 square units.
Height=x
base=x+15

Answer by jojo14344(1030)

We start with the formula, Ap=base*height -----> basic eqn.
100squnits=x*(x+15)
100=x^2+15x
x^2+15x-100=0
(x+20)(x-5)=0
2 values:
x=-20 ----------> don't use, "negative"
x=5 ------------> the one to used = height
x+15=5+15=20 ---> = base
Go back basic eqn to check,
100=5*20
100square units=100 square units
Thank you,
Jojo